OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

point-slope: (y-y1)=m(x-x1) (y--2)=2(x-3) then all you do is distribute

OpenStudy (anonymous):

*and solve for y

OpenStudy (anonymous):

Y=2x+C to find C put x=3, y=-2 then you'll have C= 8 and equation is Y=2X+8

OpenStudy (anonymous):

for slope intercept \[\frac{ Y }{ 8 }=\frac{ X }{ 4 }+1\] just divide by 8

OpenStudy (anonymous):

Thank you all.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Consider the equation 3x + 5y = 45. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points)

OpenStudy (anonymous):

np ^^

OpenStudy (anonymous):

part 1: first solve for y by subtracting 3x and then divide everything by 5.

OpenStudy (anonymous):

to graph: your equation should be y= -3/5x+9 so first graph the y-intercept by starting at 0 and go up 9 units. then use the slope by going down 3 units then to the right 5 units

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

no problem ^^